# Modeling sea ice in the marginal ice zone as a dense granular flow with rheology inferred from a discrete element model

### v1

@@ Abstract
The marginal ice zone (MIZ) represents the periphery of the sea ice cover. Here, the macroscale behavior of the sea ice results from collisions and enduring contact between ice floes. This configuration closely resembles that of dense granular flows, which have been modeled successfully with the $\mu(I)$ rheology. Here, we present a continuous model based on the $\mu(I)$ rheology which treats sea ice as a compressible fluid, with the local sea ice concentration given by a dilatancy function $\Phi(I)$. We infer expressions for $\mu(I)$ and $\Phi(I)$ from a discrete element method (DEM) which considers polygonal-shaped ice floes. We do this by driving the sea ice with a one-dimensional shearing ocean current. The resulting continuous model is a nonlinear system of equations with the sea ice velocity, local concentration, and pressure as unknowns. The rheology is given by the sum of a plastic and a viscous term. In the context of a periodic patch of ocean, which is effectively a one dimensional problem, and under steady conditions, we prove this system to be well-posed, present a numerical algorithm for solving it, and compare its solutions to those of the DEM. These comparisons demonstrate the continuous model’s ability to capture most of the DEM’s results accurately. The continuous model is particularly accurate for ocean currents faster than 0.25 m/s; however, for low concentrations and slow ocean currents, the continuous model is less effective in capturing the DEM results. In the latter case, the lack of accuracy of the continuous model is found to be accompanied by the breakdown of a balance between the average shear stress and the integrated ocean drag extracted from the DEM. Since this balance is expected to hold independently of our choice of rheology, this finding indicates that continuous models might not be able to describe sea ice dynamics for low concentrations and slow ocean currents.
@@ 1 Introduction
The periphery of the ice cover is known as the marginal ice zone (MIZ)and consists of relatively small, typically polygon-shaped ice floes. It is often defined as the region where ocean waves play an important role in shaping the morphological properties of the ice (Dumont,, 2022). On large scales, sea ice dynamics are typically described with Hibler’s model (Hibler III,, 1979), which treats ice as a viscoplastic fluid whose yield strength depends on the sea ice concentration and thickness. This model was developed for the central ice pack, where ice floes are closely interlocked and deformation is mostly due to the opening of leads or the formation of ridges.
@@
In the MIZ, however, it is the collisions and enduring contact between ice floes that give rise to the macroscale dynamical properties of the ice cover (Feltham,, 2005; Herman,, 2011, 2022). This configuration closely resembles that of dense granular flows, albeit at different spatial scales, since practically all studies for granular materials consider e.g. polystyrene beads, glass beads, and sand, whose particles’ diameters are in the order of 0.1 and 1 mm (MiDi,, 2004). Dense granular flows have been successfully modeled with the so-called $\mu(I)$ rheology (Da Cruz et al.,, 2005). The dense granular flow regime is understood as a transition between the quasi-static and dilute flow regimes. Whenever grain inertia is negligible, a quasi-static regime emerges which is often modeled as an elastoplastic solid (Nedderman,, 1992). The critical state at which plastic deformation occurs is characterized with a Coulomb-like criterion dependent on a so-called internal angle of friction (Wood,, 1990). Conversely, under great agitation and/or dilute concentrations of grains, particles interact only through binary, instantaneous, uncorrelated collisions. As a result, ideas from kinetic theory become applicable in this dilute regime (Jenkins and Savage,, 1983). However, in dense granular flows, grains interact via collisions and enduring contacts, such that inertial effects are important yet the collisions may no longer be assumed to be binary, instantaneous, or uncorrelated in general. This transitional regime is characterized in terms of the inertial number $I$ and an effective friction coefficient $\mu$ depedent on $I$ (Da Cruz et al.,, 2005).
@@
Existing models for the MIZ recognize the importance of both collisions and plastic deformation, and derive rheological models based on first principles (Shen et al.,, 1987; Gutfraind and Savage,, 1997; Feltham,, 2005). Recently, Herman, (2022) suggests the use of the $\mu(I)$ rheology for modeling sea ice in the MIZ and derives a $\mu(I)$ function from computations performed with a discrete element method (DEM). In these computations, disk-shaped ice floes are sheared by a moving wall in the classical manner of rheological studies. Unlike the previous models for the MIZ, Herman, (2022) infers the rheological properties from data generated by a DEM. In particular, Herman, (2022) fits a $\mu(I)$ function to the DEM’s data, although the resulting continuous model and its accuracy in replicating the DEM’s results is not examined.
@@
This work represents an advance in the development of a continuous model for the MIZ that could improve the accuracy of Hibler’s model, which is currently used in large-scale climate models over the MIZ, see for example (Danabasoglu et al.,, 2020). A comparison of Hibler’s model with the DEM data is presented in section 6.1, where it is demonstrated that it cannot capture the DEM results accurately. We extend the investigation initiated in Herman, (2022) and explore the $\mu(I)$ rheology’s accuracy in modeling sea ice dynamics in the MIZ. We infer a $\mu(I)$ function from computations performed with the DEM implemented in SubZero (Manucharyan and Montemuro,, 2022). This DEM considers polygon-shaped ice floes that are driven by oceanic currents in an open patch of ocean, a setup which we believe to be more natural for studying sea ice than the classical shearing test with a moving wall and disk-shaped ice floes. This inference results in a continuous viscous fluid model whose rheology is given by the sum of a viscous and a plastic term. Moreover, for this system to be well-posed, the emerging continuous model problem requires the continuum to be compressible and complemented with a constraint on the global sea ice concentration. Assuming the continuum to be compressible requires the inference of a dilatancy function $\Phi(I)$ from the DEM computations which establishes a relationship between local sea ice concentration and the inertial number $I$.
@@
The contributions of this paper can be summarized as follows: (1) Inference of the $\mu(I)$ and $\Phi(I)$ constitutive functions for sea ice in the MIZ from DEM computations performed in an open ocean configuration where the sea ice is sheared by ocean currents. (2) Analysis of the resulting continuous model, establishing the existence and uniqueness of solutions. (3) Determination of the continuous model’s range of validity by comparing its numerical solutions to those of the DEM.
@@
We remark that the analysis of the continuous model and its comparisons with the DEM are restricted to a steady one dimensional setup. The model can be extended to unsteady two dimensional problems as explained in section 2.1, although we expect new complications will arise with these extensions. For example, Barker et al., (2015) demonstrates the emergence of time-dependent instabilities in $\mu(I)$ models. These potential complications should be studied carefully in future investigations.
@@
This paper is structured as follows. In section 2, we formulate the continuous model, first in a general two-dimensional unsteady setting, then in the one dimensional steady configuration considered in this paper. In this formulation, two functions, $\mu(I)$ and $\Phi(I)$, are to be inferred from the DEM. This inference is presented in section 3. Section 4 contains a detailed analysis of the continuous model resulting from this inference. This analysis examines several properties of the momentum equation, the numerical solution of the continuous model, and its well-posedness. Then, in section 5, we compare the continuous model and the DEM. This comparison allows us to establish the range of validity of the numerical model and its limitations. In section 6, we discuss the similarities and differences between our continuous model and other sea ice models, such as Hibler’s model. We then end this paper with section 7, where we recommend potential extensions of this work to be explored in the future.
2 Mathematical formulation of the continuous problem
@@
The dense flow regime represents a transition between the quasi-static and dilute flow regimes (Da Cruz et al.,, 2005). This transitional regime is characterized in terms of the inertial number $I$, an effective friction coefficient $\mu(I)$, and, whenever the continuum is assumed to be compressible, a dilatancy function $\Phi(I)$. Below, we define these three terms and present a general formulation of the $\mu(I)$ rheology in two dimensions and in the one dimensional steady configuration considered in the subsequent sections of this paper.
@@ 2.1 The two-dimensional setting
see Hibler III, (1979) or Gray and Morland, (1994). For any scalar or vector-valued function $f$, the material derivative is given by $\mathrm{D}f/\mathrm{D}t=\partial f/\partial t+(\boldsymbol{u}\cdot\nabla)\,f$. Here, we assume the ice thickness $H$ to be spatially uniform for simplicity, although in general we require an additional equation, analogous to (2.4b) but in terms of $H$, for mass to be conserved. Equation (2.4a) is a depth-averaged statement of conservation of momentum of the sea ice layer; here, $\boldsymbol{t}_{o}$ is the drag force exerted by the ocean on the sea ice. Given the surface ocean velocity field $\boldsymbol{u}_{o}$, this drag force is generally parameterized in terms of the drag coefficient $C_{o}$ and the ocean water density $\rho_{o}$by
@@
where $\rho_{i}$ is the ice density, $\overline{d}$ an average of the ice floe size, and $p$ the pressure emerging in (2.6) (Da Cruz et al.,, 2005; Jop et al.,, 2006; Pouliquen et al.,, 2006). Throughout this document, we set $\overline{d}$ to be spatially constant over the whole domain, avoiding the need to consider the transport of this quantity. Savage, (1984) interprets the quantity $I^{2}$ as the ratio between collisional (i.e. inertial) stresses and the total shear stress exerted on the material. Accordingly, for low values of $I$, the inertial effects of grains become negligible and the flow approaches a quasi-static regime; conversely, as $I$ increases, collisional forces become increasingly important relative to the external forces exerted on the material. The functional relationship that establishes the material’s rheology is written in terms of an effective friction $\mu(I)$, defined as
@@
We remark that the effective friction $\mu$ is defined in analogy with Coulomb’s model of friction as the ratio between the shear (tangential) stress and the pressure (normal stress). Moreover, it is also helpful to think of the pressure $p$ as a quantification of the material’s strength and its resistance to viscous and plastic deformation, as made clear in section 4.
To obtain a relationship between stress and strain, we need an additional constitutive law. Jop et al., (2006) propose the following equality that aligns $\boldsymbol{\mathrm{S}}$ with $\boldsymbol{\tau}$:
@@
see Da Cruz et al., (2005). In general, $\Phi$ is found to be a strictly decreasing function of $I$, in such a way that the concentration $A$ decreases with the rate of shearing $\|\boldsymbol{\mathrm{S}}\|$, a phenomenon know as dilatancy. Moreover, if $\Phi$ is strictly decreasing, it is invertible, and one can write an expression for the pressure $p$ analogous to an equation of state in thermodynamics,
@@ 2.2 The steady one dimensional periodic ocean problem
This configuration renders the continuous problem one dimensional and independent of time, such that $\boldsymbol{u}(x,y,t)=(u(y),0)$, $A(x,y,t)=A(y)$, and $p(x,y,t)=p(y)$. In this setting, the equations for conservation of momentum (2.4a), together with the constitutive equation (2.12), simplify to the following system on $(0,L)$:
@@
Due to (2.15b), which represents the balance of momentum in the $y$-direction, the pressure is a constant (but unknown) over the domain. Da Cruz et al., (2005) and Herman, (2022) find $p$ by enforcing normal stress boundary conditions along a boundary of the domain, but we cannot do the same because the domain is periodic. In the DEM computations, which we introduce in section 3, we set a global ice concentration $A_{0}\in[0,1]$ which equals the domain averaged value of the local concentration $A$, such that
@@ 3 Inferring the constitutive equations of the system from the DEM
The system of equations presented in section 2.2 is incomplete because we need additional expressions for $\mu(I)$ and $\Phi(I)$. In this paper, we infer these additional equations from data generated with SubZero, a DEM developed by Manucharyan and Montemuro, (2022) and used for modeling sea ice dynamics with poylgonal-shaped ice floes. Following the setup presented in section 2.2, we perform runs with $n=2000$ floes over a square patch of ocean of length $L=100\,$\mathrm{k}\mathrm{m}$$, driven by the ocean velocity field (2.17).
For a given number of floes and a global sea ice concentration $A_{0}$, the initial configuration of ice floes is generated with SubZero’s packing algorithm, which is based on a Voronoi tessellation of the domain (Manucharyan and Montemuro,, 2022) (see panel (a) of figure 3 below for an example of the outcome of this packing algorithm). Defining the floe size $d$ as the square root of the floe area, this generates a polydisperse floe size distribution whose histogram we can see in Figure 2 for three values of $A_{0}$; we find that $d/L$ is approximately between 0.01 and 0.04.
@@
We run the DEM simulations for $2\times 10^{4}$ time steps, each of $5\,$\mathrm{s}$$ (the total running time is approximately equal to 27.8 hours). In general, we find that the velocity, stress, and sea ice concentration, averaged over the last 25% of the time steps, remain relatively unchanged when a longer computation is performed, and hence we consider that a steady state has been reached. This temporal averaging is performed over data which, at each time step, has been averaged spatially as described in appendix A.2. For these simulations, the material parameters which determine the effects of the ocean drag and the collisions among floes are presented in table 1. Collisional forces and the resultant stresses, which determine the fields $\boldsymbol{\sigma}$ and $p$, are computed as explained in appendix A.1.
In figure 4 we plot the values of the friction $\mu=|\sigma_{xy}|/p$ and local concentration $A$ against $I$ for different global concentrations $A_{0}$ between $0.7$ and $0.95$ and different maximum ocean velocities $u_{o,max}$ between $0.1$ and $1\,$\mathrm{m}\mathrm{/}\mathrm{s}$$. The mean ocean velocities in the ocean patch are therefore between $0.05$ and 0.5 m/s, values that are consistent with real observations (Stewart et al.,, 2019). In all of these computations, we set the ice thickness to $H=2\,$\mathrm{m}$$. We find an increase in the friction $\mu$ and a decrease in the local concentration $A$ as $I$ increases. The decrease in the local concentration of sea ice $A$ with an increase in $I$ is due to dilatancy. Figure 3 presents an example of how dilatancy emerges in the DEM computations: given a random initial distribution of ice floes, when a steady state is reached the concentration $A$ decreases in the areas where the largest shearing occurs ($y=1/4$ and $y=3/4$), and increases elsewhere. Since the global sea ice concentration $A_{0}$ is constant, in this context dilatancy represents a reorganization of the local concentration profile $A(y)$.
@@
A linear behaviour is also found for $\mu$ in Da Cruz et al., (2005). The four parameters $(\mu_{0},\mu_{1},\phi_{0},\alpha)$ are calculated by minimizing the least-squares misfit problem between the points in figure 4 and the functions in (3.1). The resulting values are shown in table 2. For the remainder of the document, any numerical solution of the one dimensional system (2.15) is solved by setting the parameters in functions (3.1) to the values given in table 2.
@@
Table 2: Parameters for the functions $\mu(I)$ and $\Phi(I)$ in (3.1) obtained by minimizing the least-squares misfit with the data plotted in figure 4. These are the numerical values used for computing solutions to the continuous model (2.15) throughout this paper.
Departure from the fitting curves are most visible when the ocean velocities and sea ice concentrations are low, see the case where $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}=0.7$. Unsurprisingly, in section 5 below, we also find the greatest misfit between the DEM and the continuous model precisely in this setting, when $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}=0.7$, see panel (m) in figure 9 below. In particular, in this setting, the fundamental balance between shear stress and ocean drag in the DEM is found to no longer hold, see section 5.
@@
A consequence of this linear behavior is that $\mu$ is approximately constant for small values of $I$, and this is precisely what we see for $I<10^{-2}$ in figure 4. For $I\ll 1$, we have that $\mu(I)\approx\mu_{0}$, and therefore it is the plastic term that dominates the rheology. This is essentially the quasi-static regime, where collisions are negligible. This plastic term follows from a Mohr-Coulomb yield criterion with an internal angle of friction $\tan^{-1}(\mu_{0})$; examples of the Mohr-Coulomb yield criterion used for sea ice modeling can be found in Ip et al., (1991), Gutfraind and Savage, (1997) and Ringeisen et al., (2019). The viscous term, which becomes increasingly important as the inertial number $I$ increases, can be associated with the collisional component of the rheology. A viscous rheology is derived in Shen et al., (1987) for modeling the rheological effects of collisions in sea ice, which, as we explain in section 6, is very similar to the viscous component in (3.2).

### v2

# Modeling sea ice in the marginal ice zone as a dense granular flow with rheology inferred from discrete element model data

@@ Abstract
The marginal ice zone (MIZ) represents the periphery of the sea ice cover. In this region, the macroscale behavior of the sea ice results from collisions and enduring contact between ice floes. This configuration closely resembles that of dense granular flows, which have been modeled successfully with the $\mu(I)$ rheology. Here, we present a continuum model based on the $\mu(I)$ rheology which treats sea ice as a compressible fluid, with the local sea ice concentration given by a dilatancy function $\Phi(I)$. We infer expressions for $\mu(I)$ and $\Phi(I)$ by nonlinear regression using data produced with a discrete element method (DEM) which considers polygonal-shaped ice floes. We do this by driving the sea ice with a one-dimensional shearing ocean current. The resulting continuum model is a nonlinear system of equations with the sea ice velocity, local concentration, and pressure as unknowns. The rheology is given by the sum of a plastic and a viscous term. In the context of a periodic patch of ocean, which is effectively a one dimensional problem, and under steady conditions, we prove this system to be well-posed, present a numerical algorithm for solving it, and compare its solutions to those of the DEM. These comparisons demonstrate the continuum model’s ability to capture most of the DEM’s results accurately. The continuum model is particularly accurate for ocean currents faster than 0.25 m/s; however, for low concentrations and slow ocean currents, the continuum model is less effective in capturing the DEM results. In the latter case, the lack of accuracy of the continuum model is found to be accompanied by the breakdown of a balance between the average shear stress and the integrated ocean drag extracted from the DEM. Since this balance is expected to hold independently of our choice of rheology, this finding indicates that continuum models might not be able to describe sea ice dynamics for low concentrations and slow ocean currents.
@@ 1 Introduction
The periphery of the ice cover is known as the marginal ice zone (MIZ) and consists of relatively small, typically polygon-shaped ice floes. It is often defined as the region where ocean waves play an important role in shaping the morphological properties of the ice (Dumont, 2022). On large scales, sea ice dynamics are typically described with Hibler’s model (Hibler III, 1979), which treats ice as a viscoplastic fluid whose yield strength depends on the sea ice concentration and thickness. This model was developed for the central ice pack, where ice floes are closely interlocked and deformation is mostly due to the opening of leads or the formation of ridges. Recently, elasto-brittle rheologies have also been used to model the evolution of the central ice pack. This class of rheological models, which were first proposed by Girard et al. (2011), appear to be superior to Hibler’s model in capturing the ice deformation field (Rampal et al., 2019).
@@
In the MIZ, however, it is the collisions and enduring contact between ice floes that give rise to the macroscale dynamical properties of the ice cover (Feltham, 2005; Herman, 2011, 2022). This configuration closely resembles that of dense granular flows, albeit at different spatial scales, since practically all studies for granular materials consider e.g. polystyrene beads, glass beads, and sand, whose particles’ diameters are in the order of 0.1 and 1 mm (MiDi, 2004). Dense granular flows have been successfully modeled with the so-called $\mu(I)$ rheology (Da Cruz et al., 2005). The dense granular flow regime is understood as a transition between the quasi-static and dilute flow regimes. Whenever grain inertia is negligible, a quasi-static regime emerges which is often modeled as an elastoplastic solid (Nedderman, 1992). The critical state at which plastic deformation occurs is characterized with a Coulomb-like criterion dependent on a so-called internal angle of friction (Wood, 1990). Conversely, under great agitation and/or dilute concentrations of grains, particles interact only through binary, instantaneous, uncorrelated collisions. As a result, ideas from kinetic theory become applicable in this dilute regime (Jenkins & Savage, 1983). However, in dense granular flows, grains interact via collisions and enduring contacts, such that inertial effects are important yet the collisions may no longer be assumed to be binary, instantaneous, or uncorrelated in general. This transitional regime is characterized in terms of the inertial number $I$ and an effective friction coefficient $\mu$ dependent on $I$ (Da Cruz et al., 2005).
@@
Existing models for the MIZ recognize the importance of both collisions and plastic deformation, and derive rheological models based on first principles (Shen et al., 1987; Gutfraind & Savage, 1997; Feltham, 2005). Recently, Herman (2022) suggests the use of the $\mu(I)$ rheology for modeling sea ice in the MIZ and derives a $\mu(I)$ function from computations performed with a discrete element method (DEM). In these computations, disk-shaped ice floes are sheared by a moving wall in the classical manner of rheological studies. Unlike the previous models for the MIZ, Herman (2022) infers the rheological properties from data generated by a DEM. In particular, Herman (2022) fits a $\mu(I)$ function to the DEM’s data, although the resulting continuum model and its accuracy in replicating the DEM’s results is not examined.
@@
This work represents an advance in the development of a continuum model for the MIZ that could improve the accuracy of Hibler’s model, which is currently used in large-scale climate models over the MIZ, see for example (Danabasoglu et al., 2020). A comparison of Hibler’s model with the DEM data is presented in section 6.1, where it is demonstrated that it cannot capture the DEM results accurately. We extend the investigation initiated in Herman (2022) and explore the $\mu(I)$ rheology’s accuracy in modeling sea ice dynamics in the MIZ. We infer a $\mu(I)$ function from data produced with the DEM implemented in SubZero (Manucharyan & Montemuro, 2022). This DEM considers polygon-shaped ice floes that are driven by oceanic currents in an open patch of ocean, a setup which we believe to be more natural for studying sea ice than the classical shearing test with a moving wall and disk-shaped ice floes. This inference results in a continuum viscous fluid model whose rheology is given by the sum of a viscous and a plastic term. Moreover, for this system to be well-posed, the emerging model problem requires the continuum to be compressible and complemented with a constraint on the global sea ice concentration. Assuming the continuum to be compressible requires the inference of a dilatancy function $\Phi(I)$ from the DEM computations which establishes a relationship between local sea ice concentration and the inertial number $I$.
@@
The contributions of this paper can be summarized as follows: (1) Inference of the $\mu(I)$ and $\Phi(I)$ constitutive functions for sea ice in the MIZ from data produced with the DEM. These computations are performed in an open ocean configuration where the sea ice is sheared by ocean currents. (2) Analysis of the resulting continuum model, establishing the existence and uniqueness of solutions. (3) Determination of the continuum model’s range of validity by comparing its numerical solutions to those of the DEM.
@@
We remark that the analysis of the continuum model and its comparisons with the DEM are restricted to a steady one dimensional setup. The model can be extended to unsteady two dimensional problems as explained in section 2.1, although we expect new complications will arise with these extensions. For example, Barker et al. (2015) demonstrate the emergence of time-dependent instabilities in $\mu(I)$ models, which Schaeffer et al. (2019) remedy with further modifications of the model. These potential complications should be studied carefully in future investigations.
@@
This paper is structured as follows. In section 2, we formulate the continuum model, first in a general two-dimensional unsteady setting, then in the one dimensional steady configuration considered in this paper. In this formulation, two functions, $\mu(I)$ and $\Phi(I)$, are to be inferred from DEM data. This inference is presented in section 3. Section 4 contains a detailed analysis of the continuum model resulting from this inference. This analysis examines several properties of the momentum equation, the numerical solution of the continuum model, and its well-posedness. Then, in section 5, we compare the continuum model and the DEM. This comparison allows us to establish the range of validity of the continuum model and its limitations. In section 6, we discuss the similarities and differences between our continuum model and other sea ice models, such as Hibler’s model. We then end this paper with section 7, where we recommend potential extensions of this work to be explored in the future.
2 Mathematical formulation of the continuum problem
@@
The dense flow regime represents a transition between the quasi-static and dilute flow regimes (Da Cruz et al., 2005). This transitional regime is characterized in terms of the inertial number $I$, an effective friction coefficient $\mu(I)$, and, whenever the continuum is assumed to be compressible, a dilatancy function $\Phi(I)$. Below, we define these three terms and present a general formulation of the $\mu(I)$ rheology in two dimensions and in the one dimensional steady configuration considered in the subsequent sections of this paper.
@@ 2.1 The two-dimensional setting
see Hibler III (1979) or Gray & Morland (1994). For any scalar or vector-valued function $f$, the material derivative is given by $\mathrm{D}f/\mathrm{D}t=\partial f/\partial t+(\boldsymbol{u}\cdot\nabla)\,f$. Here, we assume the ice thickness $H$ to be spatially uniform for simplicity, although in general we require an additional equation, analogous to (2.4b) but in terms of $H$, for mass to be conserved. Equation (2.4a) is a depth-averaged statement of conservation of momentum of the sea ice layer; here, $\boldsymbol{t}_{o}$ is the drag force exerted by the ocean on the sea ice. Given the surface ocean velocity field $\boldsymbol{u}_{o}$, this drag force is generally parameterized in terms of the drag coefficient $C_{o}$ and the ocean water density $\rho_{o}$ by
@@
where $\rho_{i}$ is the ice density, $\overline{d}$ denotes an average ice floe size, and $p$ is the pressure emerging in (2.6) (Da Cruz et al., 2005; Jop et al., 2006; Pouliquen et al., 2006). Throughout this document, we set $\overline{d}$ to be spatially constant over the whole domain, avoiding the need to consider the transport of this quantity. Savage (1984) interprets the quantity $I^{2}$ as the ratio between collisional (i.e. inertial) stresses and the total shear stress exerted on the material. Accordingly, for low values of $I$, the inertial effects of grains become negligible and the flow approaches a quasi-static regime; conversely, as $I$ increases, collisional forces become increasingly important relative to the external forces exerted on the material. The functional relationship that establishes the material’s rheology is written in terms of an effective friction $\mu(I)$, defined as
@@
We remark that the effective friction $\mu$ is defined in analogy with Coulomb’s model of friction as the ratio between the shear (tangential) stress and the pressure (normal stress). Moreover, it is also helpful to think of the pressure $p$ as a quantification of the material’s strength and its resistance to viscous and plastic deformation, as made clear in section 4. It should be noted that the $\mu(I)$ model is a phenomenological model that has been found to work well with granular media, yet it is unclear if it represents some kind of limit for a large particle system.
To obtain a relationship between stress and strain, we need an additional constitutive law. Jop et al. (2006) propose the following equality that aligns $\boldsymbol{\mathrm{S}}$ with $\boldsymbol{\tau}$:
@@
see Da Cruz et al. (2005). In general, $\Phi$ is found to be a strictly decreasing function of $I$, in such a way that the concentration $A$ decreases with the rate of shearing $\|\boldsymbol{\mathrm{S}}\|$, a phenomenon know as dilatancy. Moreover, if $\Phi$ is strictly decreasing, it is invertible, and one can write an expression for the pressure $p$ analogous to an equation of state in thermodynamics,
@@ 2.2 The steady one dimensional periodic ocean problem
This configuration renders the continuum problem one dimensional and independent of time, such that $\boldsymbol{u}(x,y,t)=(u(y),0)$, $A(x,y,t)=A(y)$, and $p(x,y,t)=p(y)$. In this setting, the equations for conservation of momentum (2.4a), together with the constitutive equation (2.12), simplify to the following system on $(0,L)$:
@@
Due to (2.15b), which represents the balance of momentum in the $y$-direction, the pressure is a constant (but unknown) over the domain. Da Cruz et al. (2005) and Herman (2022) find $p$ by enforcing normal stress boundary conditions along a boundary of the domain, but we cannot do the same because the domain is periodic. In the DEM computations, which we introduce in section 3, we set a global ice concentration $A_{0}\in[0,1]$ which equals the domain averaged value of the local concentration $A$, such that
@@ 3 Inferring the constitutive equations of the system from the DEM
The system of equations presented in section 2.2 is incomplete because we need additional expressions for $\mu(I)$ and $\Phi(I)$. In this paper, we infer these additional equations from data generated with SubZero, a DEM developed by Manucharyan & Montemuro (2022) and used for modeling sea ice dynamics with polygonal-shaped ice floes. Following the setup presented in section 2.2, we perform runs with $n=2000$ floes over a square patch of ocean of length $L=100\,$\mathrm{k}\mathrm{m}$$, driven by the ocean velocity field (2.17). This means that the $\mu(I)$ and $\Phi(I)$ functions are inferred from DEM solutions with a constant number of floes (and therefore constant average floe size). In theory, the effects of floe size are included in the non-dimensional parameter $I$. In practice, the effects of $n$ are subtle and not too well captured with our continuum model, see figure 12 in section 5 below.
For a given number of floes and a global sea ice concentration $A_{0}$, the initial configuration of ice floes is generated with SubZero’s packing algorithm, which is based on a Voronoi tessellation of the domain (Manucharyan & Montemuro, 2022) (see panel (a) of figure 2 below for an example of the outcome of this packing algorithm). Defining the floe size $d$ as the square root of the floe area, this generates a polydisperse floe size distribution whose histogram we can see in figure 3 for three values of $A_{0}$; we find that $d/L$ is approximately between 0.01 and 0.04. Although we do not study the effects of polydispersity on the rheology here, we remark that Herman (2022) finds the dilatancy law $\Phi$ to vary visibly with the degree of polydispersity, while the effective friction $\mu$ presents much smaller variations.
We run the DEM simulations for $2\times 10^{4}$ time steps, each of $5\,$\mathrm{s}$$ (the total running time is approximately equal to 27.8 hours). In general, we find that the velocity, stress, and sea ice concentration, averaged over the last 25% of the time steps, remain relatively unchanged when a longer computation is performed, and hence we consider that a steady state has been reached. This temporal averaging is performed over data which, at each time step, has been averaged spatially over a grid as described in appendix A.2. For these simulations, the material parameters which determine the effects of the ocean drag and the collisions among floes are presented in table 1. Collisional forces and the resultant stresses, which determine the fields $\boldsymbol{\sigma}$ and $p$, are computed as explained in appendix A.1.
In figure 4 we plot the values of the friction $\mu=|\sigma_{xy}|/p$ and local concentration $A$ against $I$ for different global concentrations $A_{0}$ between $0.7$ and $0.95$ and different maximum ocean velocities $u_{o,max}$ between $0.1$ and $1\,$\mathrm{m}\mathrm{/}\mathrm{s}$$. The mean ocean velocities in the ocean patch are therefore between $0.05$ and 0.5 m/s, values that are consistent with real observations (Stewart et al., 2019). In all of these computations, we set the ice thickness to $H=2\,$\mathrm{m}$$. We find an increase in the friction $\mu$ and a decrease in the local concentration $A$ as $I$ increases. The decrease in the local concentration of sea ice $A$ with an increase in $I$ is due to dilatancy. Figure 2 presents an example of how dilatancy emerges in the DEM computations: given a random initial distribution of ice floes, when a steady state is reached the concentration $A$ decreases in the areas where the largest shearing occurs ($y=1/4$ and $y=3/4$), and increases elsewhere. Since the global sea ice concentration $A_{0}$ is constant, in this context dilatancy represents a reorganization of the local concentration profile $A(y)$.
@@
A linear behavior is also found for $\mu$ in Da Cruz et al. (2005). The four parameters $(\mu_{0},\mu_{1},\phi_{0},\alpha)$ are calculated by minimizing the least-squares misfit problem between the points in figure 4 and the functions in (3.1). The resulting values are shown in table 2. For the remainder of the document, any numerical solution of the one dimensional system (2.15) is solved by setting the parameters in functions (3.1) to the values given in table 2.
@@
Table 2: Parameters for the functions $\mu(I)$ and $\Phi(I)$ in (3.1) obtained by minimizing the least-squares misfit with the data plotted in figure 4. These are the numerical values used for computing solutions to the continuum model (2.15) throughout this paper.
Departure from the fitting curves are most visible when the ocean velocities and sea ice concentrations are low, see the case where $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}=0.7$. Unsurprisingly, in section 5 below, we also find the greatest misfit between the DEM and the continuum model precisely in this setting, when $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}=0.7$, see panel (m) in figure 9 below. In particular, in this setting, the fundamental balance between shear stress and ocean drag in the DEM is found to no longer hold, see section 5.
@@
A consequence of this linear behavior is that $\mu$ is approximately constant for small values of $I$, and this is precisely what we see for $I<10^{-2}$ in figure 4. For $I\ll 1$, we have that $\mu(I)\approx\mu_{0}$, and therefore it is the plastic term that dominates the rheology. This is essentially the quasi-static regime, where collisions are negligible. This plastic term follows from a Mohr-Coulomb yield criterion with an internal angle of friction $\tan^{-1}(\mu_{0})$; examples of the Mohr-Coulomb yield criterion used for sea ice modeling can be found in Ip et al. (1991), Gutfraind & Savage (1997) and Ringeisen et al. (2019). The viscous term, which becomes increasingly important as the inertial number $I$ increases, can be associated with the collisional component of the rheology. A viscous rheology is derived in Shen et al. (1987) for modeling the rheological effects of collisions in sea ice, which, as we explain in section 6, is very similar to the viscous component in (3.2).